# Robust Predictive Motion Planning by Learning Obstacle Uncertainty

### v1

@@ VI Simulations of Reach-Avoid Motion Planning
The three methods are further compared through $100$ Monte-Carlo simulations with a randomly sampled initial state for the SV, while other conditions are the same as in Fig. 7. The planners are compared by counting the collision-free rate of planning, where a collision-free mission is fulfilled if the EV has no collision with either the SV or the driving-area boundary. Among all successful cases, the completed cases are counted, where a complete mission means that the Euclidean distance between the state of the EV and its reference state converges to less than or equal to $0.2$ within a predefined time limitation ($13.75\ {\rm s}$ in the simulation). For all completed cases with each planner, the minimum polytope distance between the EV and SV of each simulation (${\rm min}(d^{e,s})$), the first time instant when the EV reaches the reference state ($\tau^{e,\rm ref}$), and the sum of the cost function value are compared, and the results are presented in Table III. It is seen that the proposed method and RMPC achieve a collision-free rate of $100\%$, whereas DMPC achieves $76\%$. Among all collision-free cases, the proposed method and DMPC have a $100\%$ complete rate, while the complete rate for RMPC is $73\%$. Analyzing the random variables ${\rm min}(d^{e,s})$, $\tau^{e,\rm ref}$, and the sum of cost reveals that both the proposed method and DMPC guide the EV to the reference state at the same time instant, and the costs are close. However, the proposed method demonstrates superior safety compared to DMPC. On the other hand, RMPC ensures a large distance between the EV and SV in each simulation, but it takes more time and a larger cost to converge.
The average execution time for each iteration of the proposed method, DMPC, and RMPC is $0.105\ {\rm s}$, $0.056\ {\rm s}$, and $0.103\ {\rm s}$, respectively, with standard deviations of $0.025\ {\rm s}$, $0.023\ {\rm s}$, and $0.022\ {\rm s}$, respectively. These results were recorded from $10$ random simulations where each MPC algorithm was run $55$ steps. The simulation results demonstrate that the proposed method based on learning obstacle uncertainties performs successfully in the absence of prior knowledge of the environment and the obstacle. It is as safe as the RMPC, while simultaneously enhancing efficiency and reducing the cost associated with motion-planning tasks. In contrast, when compared to the DMPC, it markedly increases safety without incurring additional costs or compromising the efficiency of executing the tasks.
@@
The DMPC and RMPC are selected as the baseline methods for two primary reasons. First, in the studied reach-avoid problem, the DMPC excels in planning the most efficient reference trajectory when the motion-planning problem is feasible, while the RMPC is known for its guaranteed safety. The evaluations of the proposed method are demonstrated by comparing it with the efficiency-optimal method and the safety-optimal method, respectively. Second, similar to the proposed method, the DMPC and RMPC rely on a few assumptions on the traffic scenario. Other methods such as risk-aware optimal control [4], interaction and safety-aware MPC [24], and non-conservative stochastic MPC [23], [27], require certain assumptions on uncertainty distributions of the SVs, which are not satisfied in our case studies. Therefore, DMPC and RMPC provide a fair comparison for assessing the proposed method’s performance.
Remark VI.2.
In the implementations, one important step is to, according to Assumption IV.1, estimate the control input $u_{t-1}^{s}$ of model (4) of the SV at every time step $t$. This is achieved by observing the velocities of the SV at time steps $t-1$ and $t$ to calculate its acceleration at time step $t-1$ in the vehicle frame, which is combined with its observed orientation at time step $t-1$ to estimate $u_{t-1}^{s}$.
Remark VI.3.
The case studies consider the motion-planning problem for the EV in the presence of one SV, while the method can also manage multiple obstacles. The effectiveness of the proposed method in scenarios involving multiple obstacles has been verified in simulation, and the results are provided together with the videos in Section VII-B.
@@ VII Hardware Experiments
TABLE IV: Experimental Parameters in Each Scenario
@@
This section presents hardware-experiment results in scenarios different from those in Section VI to demonstrate the method’s applicability across different situations, as well as the real-time performance of the method in applications. The experiment videos are accessible online [^3] https://youtu.be/r8YCDAh7Hts.
@@ VII-A Scenario Description
The design of the MPC planner follows the same approach as presented in Section VI, so details are omitted for conciseness. Note that to enhance the efficiency of solving the OCP problem, several adaptations are applied. First, the dimension of the EV model in (III.1) is reduced to $4$ by taking the longitudinal acceleration and front tire angle as the inputs. Second, the prediction horizon is changed to $N=9$.
@@ VII-B Experiment Results and Discussion
In the experiments, the polytope computation and the optimization problem are solved by the same tools as in Section VI. The positioning of both EV and SV was performed by the Qualisys motion-capture system [39], and the control inputs of the vehicles were computed on a laptop running Ubuntu 20.04 LTS operating system. The communications were established by the Robot Operating System (ROS Noetic). The reference paths of the EV and SV were followed by a pure-pursuit controller [40], and the reference velocities were followed by the embedded velocity-tracking controllers of the hamster platform and the cascar platform, respectively. The initial states and reference velocities of the EV and SV in each scenario are collected in Table IV. The reference velocity for the hamster platform is designed following the specifications detailed in [41] and [42].
In the encounter scenario, the EV and SV will intersect twice. The trajectories of both vehicles during these encounters are depicted in Fig. 10. Fig. 11(a) visualizes the learned control set $\hat{\mathbb{U}}_{t}^{s}$ of the SV over time, and Fig. 11(b) presents the velocity of the SV in the ground coordinate system. Following Fig. 11(a), Fig. 12 illustrates how the predicted occupancy of the SV over the prediction horizon, i.e., $\hat{\mathcal{O}}_{i|t}^{s}$, evolves at the initial moments. Leveraging the occupancy $\hat{\mathcal{O}}_{i|t}^{s}$, the EV plans its reference by solving problem (V). The occupancy $\hat{\mathcal{O}}_{i|t}^{s}$, along with the planned reference path of the EV at key moments during the first encounter, are exhibited in Fig. 13.
@@
The trajectories of the EV in the overtaking and evasion experiments are illustrated in Fig.14 and Fig.15, respectively. In both scenarios, the proposed method successfully addresses the motion-planning challenges. The corresponding learned sets $\hat{\mathbb{U}}_{t}^{s}$ and the predicted occupancies $\hat{\mathcal{O}}_{i|t}^{s}$, along with the planned collision-avoidance trajectories of the EV, exhibit qualitative similarities. In the presented experiments, the average execution time of the proposed method is $0.064\ {\rm s}$, and the standard deviation is $0.014\ {\rm s}$, with the motion planner executed $200$ times in the encounter scenario. The computation time in the overtaking and evasion scenarios is comparably similar. Both the simulations and experiments demonstrate the method’s effectiveness in addressing environmental uncertainties and provide validation its real-time applicability.
VIII Conclusion and Future Work
@@
Future research will focus on extending the proposed method for nonlinear obstacle models and other environmental uncertainties. Another research direction of great interest is the interaction-aware motion planning of multi-agent systems by proactively learning the decision-making of other agents.
@@ Acknowledgments
The authors would like to acknowledge Prof. Lars Nielsen at Linköping University for discussions regarding the second-order integrator and polytope-based collision-avoidance constraints. Lic. Theodor Westny at Linköping University is acknowledged for skillful assistance in the experiments.

### v2

@@ VI Simulations of Reach-Avoid Motion Planning
The three methods are further compared through $300$ Monte-Carlo simulations with a randomly sampled initial state for the SV, while other conditions are the same as in Fig. 7. The planners are implemented with the prediction horizons $N=10$ and $N=8$, respectively, and are compared by counting the collision-free rate of planning, where a collision-free mission is fulfilled if the EV has no collision with either the SV or the driving-area boundary. Among all successful cases, the complete cases are counted, where a complete mission means that the Euclidean distance between the state of the EV and its reference state converges to less than or equal to $0.2$ within a predefined time limitation ($13.75\ {\rm s}$ in the simulation). For all completed cases with each planner, the minimum polytope distance between the EV and SV of each simulation ($d_{\rm min}^{e,s}$), the first time instant when the EV reaches the reference state ($\tau^{e,\rm ref}$), and the sum of the cost function value ($J_{\rm sum}$) are compared. The results are presented in Table III. It is seen that only the proposed method achieves a collision-free rate of $100\%$ for both prediction horizons. Note that the RMPC does not achieve $100\%$ with $N=10$ because the obstacle occupancy expands with the prediction horizon. This reduces the driveable space, i.e., the feasible region, for the EV in the limited area, thereby causing collisions with the boundary in some cases. Among all collision-free cases, the proposed method and DMPC have a $100\%$ complete rate, while the RMPC does not fulfill all the tasks. Analyzing the random variables $d_{\rm min}^{e,s}$, $\tau^{e,\rm ref}$, and $J_{\rm sum}$ reveals that both the proposed method and DMPC guide the EV to the reference state at the same time instant, and the costs are close. However, the proposed method demonstrates superior safety compared to DMPC. On the other hand, RMPC ensures a large distance between the EV and SV in each simulation, but it takes more time and a larger cost to converge. For the proposed method, although the prediction horizons $N=10$ and $N=8$ both result in $100\%$ collision-free rate and complete rate, the shorter horizon can cause a smaller minimum distance between the EV and SV, and a larger cost of the maneuver. The simulation results demonstrate that the proposed method performs successfully in the absence of prior knowledge, e.g., the unknown dynamics, unknown distribution of control actions, and uncertain intentions of the obstacle. Empirically, the results show that the proposed approach is safe in all considered instances, while simultaneously enhancing efficiency and feasibility and reducing the cost associated with motion-planning tasks compared with the RMPC. In contrast, when compared to the DMPC, it markedly increases safety without incurring additional costs or compromising the efficiency of executing the tasks.
The computation time for each iteration of the proposed method, DMPC, and RMPC with different prediction horizons is summarized in Table IV. The results were recorded from $20$ random simulations where each MPC algorithm was run $55$ steps. It is seen in Table IV that DMPC is the most computationally efficient, while the prediction horizon has an obvious influence on the computational efficiency of the proposed method and RMPC. In addition, for the proposed approach, the majority of the computation load comes from predicting the forward reachable set (14) and solving the MPC problem (16). The computation time for solving  (16) increases slightly with a longer prediction horizon, while the polytope-based computation of occupancy sets using  (14) can be computationally expensive when the prediction horizon is long. Table III shows that the computation time is sufficient for running the planner at around ${\rm{10\ Hz}}$ in simulations, while it further indicates that a trade-off between the performance and the computational efficiency should be carefully considered when defining the prediction horizon in practical applications.
@@
The DMPC and RMPC are selected as the baseline methods for two primary reasons. First, in the studied reach-avoid problem, the DMPC excels in planning the most efficient reference trajectory when the motion-planning problem is feasible, while the RMPC is known for its guaranteed safety when the problem is feasible. The evaluations of the proposed method are demonstrated by comparing it with the efficiency-optimal method and the safety-optimal method, respectively. Second, similar to the proposed method, the DMPC and RMPC rely on a few assumptions on the traffic scenario. Other methods such as risk-aware optimal control [4], interaction and safety-aware MPC [25], and non-conservative stochastic MPC [24], [28], require certain assumptions on uncertainty distributions of the SVs, which are not satisfied in our case studies. Therefore, DMPC and RMPC provide a fair comparison for assessing the proposed method’s performance.
@@ VII Hardware Experiments
TABLE V: Experimental Parameters in Each Scenario
@@
This section presents hardware-experiment results in scenarios different from those in Section VI to demonstrate the method’s applicability across different situations, as well as the real-time performance of the method in applications. The experiment videos are accessible online [^3] https://youtu.be/ulBvjGxYkIE.
@@ VII-A Scenario Description
The design of the MPC planner follows the same approach as presented in Section VI, so details are omitted for conciseness. Note that to enhance the efficiency of solving the OCP problem, several adaptations are applied. First, the dimension of the EV model in (3) is reduced to $4$ by taking the longitudinal acceleration and front tire angle as the inputs. Second, the prediction horizon is changed to $N=9$.
@@ VII-B Experiment Results and Discussion
In the experiments, the polytope computation and the optimization problem are solved by the same tools as in Section VI. The positioning of both EV and SV was performed by the Qualisys motion-capture system [41]. The estimation uncertainties of elements in the information set (5) arose from the state-measurement uncertainties of the SV. However, these measurement uncertainties were minor as a result of the high accuracy of the positioning system. The control inputs of the vehicles were computed on a laptop running Ubuntu 20.04 LTS operating system. The communications were established by the Robot Operating System (ROS Noetic). The reference paths of the EV and SV were followed by a pure-pursuit controller [42], and the reference velocities were followed by the embedded velocity-tracking controllers of the hamster platform and the cascar platform, respectively. The initial states and reference velocities of the EV and SV in each scenario are collected in Table V. The reference velocity for the hamster platform is designed following the specifications detailed in [43] and [44]. In the experiments, the SV was designed to execute non-smooth trajectories by oscillating the front wheels to increase motion uncertainties for the EV. The set $\hat{\mathbb{U}}_{t}^{s}$ was learned by estimating the accelerations of the SV in the ground coordinate system, and this does not require oscillation of the SV input.
In the encounter scenario, the EV and SV will intersect twice. The trajectories of both vehicles during these encounters are depicted in Fig. 10. Fig. 11(a) visualizes the learned control set $\hat{\mathbb{U}}_{t}^{s}$ of the SV over time, and Fig. 11(b) presents the velocity of the SV in the ground coordinate system. Following Fig. 11(a), Fig. 12 illustrates how the predicted occupancy of the SV over the prediction horizon, i.e., $\hat{\mathcal{O}}_{i|t}^{s}$, evolves at the initial moments. Leveraging the occupancy $\hat{\mathcal{O}}_{i|t}^{s}$, the EV plans its reference by solving problem (15). The occupancy $\hat{\mathcal{O}}_{i|t}^{s}$, along with the planned reference path of the EV at key moments during the first encounter, are exhibited in Fig. 13.
@@
The trajectories of the EV in the overtaking and evasion experiments are illustrated in Fig.14 and Fig.15, respectively. In both scenarios, the proposed method successfully addresses the motion-planning challenges. The corresponding learned sets $\hat{\mathbb{U}}_{t}^{s}$ and the predicted occupancies $\hat{\mathcal{O}}_{i|t}^{s}$, along with the planned collision-avoidance trajectories of the EV, exhibit qualitative similarities. In the presented experiments, the average execution time of the proposed method is $0.064\ {\rm s}$, and the standard deviation is $0.014\ {\rm s}$, with the motion planner executed $200$ times in the encounter scenario. The computation time in the overtaking and evasion scenarios is comparably similar. Both the simulations and experiments demonstrate the method’s effectiveness in addressing environmental uncertainties and provide validation its real-time applicability. It is also crucial to note that in both Sections VIVII, the SV is predicted by the double-integrator model (4) with the true dynamics completely unknown to the EV, this demonstrated the effectiveness of the method in the presence of unknown dynamics of obstacles.
VII-C Discussion on Applicability and Future Research
The case studies used one SV to illustrate the implementation of the proposed method, while the proposed method can directly be applied to manage more obstacles. In addition, the method applies to more scenarios, like highway overtaking, merging, and roundabout driving scenarios. The method’s effectiveness has been evaluated in four additional scenarios, including a multiple obstacles scenario, a high-speed overtaking scenario, a cornered track scenario, and a roundabout traffic scenario. The results are presented as supplementary material in the video linked in Section VII-B. Furthermore, we remark that the proposed method does not have assumptions about the obstacle’s control strategy, this means the method also handles the case where the obstacle actively reacts to the maneuver of the EV. This has been verified in an oncoming traffic scenario as in [45, Section VI. A], where both the EV and obstacle successfully generate collision-free trajectories when they apply the same strategy. The results are also provided as supplementary material in the video. Note that in extreme high-speed high-curvature scenarios, the effectiveness of this method might be affected because of the unpredictable road geometry that can alter the intended trajectories of the obstacles. These factors are not included in this paper.
Future research would focus on integrating with state-of-the-art environment-aware motion-prediction methods, e.g., the map-informed approach [46], to address problems in more challenging high-speed high-curvature scenarios. Another research direction of great interest is integrating an interaction-aware motion-prediction method, e.g., from a game theoretical perspective [47], with the proposed approach for proactive motion planning in multi-vehicle interaction scenarios.
VIII Conclusion
@@ Acknowledgments
The authors would like to acknowledge Prof. Lars Nielsen at Linköping University for discussions regarding the second-order integrator and polytope-based collision-avoidance constraints. Dr. Theodor Westny at Linköping University is acknowledged for skillful assistance in the experiments.